OpenStudy (anonymous):

Which of the following equations is written in standard form? A.18=-2y+6x B.2x+8y-1 C.4y-1=0 D.-4=2/3x-3/4y @Notamathgenius

OpenStudy (anonymous):

i think D

OpenStudy (notamathgenius):

Yea I was thinking D too. But It May be A If I am not mistaken.

OpenStudy (anonymous):

I think it;s A, because it's in y=mx+b equation

OpenStudy (notamathgenius):

^^^^ True.

OpenStudy (anonymous):

\[\frac{ x }{ a }+\frac{ y }{ b }=c\]

OpenStudy (anonymous):

or it could also be ^^

OpenStudy (anonymous):

wow we are probably making @AnastasiaSttl confused...lol

OpenStudy (notamathgenius):

Ok its Either A or D we must come to a conclusive agreement.

OpenStudy (anonymous):

I think its A

OpenStudy (notamathgenius):

thats what I was gonna say but I dont want to start anything.

OpenStudy (anonymous):

me too:) 2/4 @Notamathgenius ?? @lovemenplayhard ??

OpenStudy (notamathgenius):

I would go with A @AnastasiaSttl

OpenStudy (anonymous):

I just put A It was correct, Can you help me with the new question I asked & thank you !!

OpenStudy (anonymous):

shoot

OpenStudy (notamathgenius):

Yea we are smart :D

OpenStudy (anonymous):

yea